OpenStudy (anonymous):

\[3m-9\m^2-2m \div m^2-m-6\m^2-4\]

OpenStudy (anonymous):

3m-9/m^2-2m divided by m^2-m-6/m^2-4

OpenStudy (anonymous):

sorry i cant get it to type correct

OpenStudy (mertsj):

\[\frac{3(m-3)}{m(m-2)}\times \frac{(m-2)(m+2)}{(m-3)(m+2)}=\frac{3}{m}\]

OpenStudy (anonymous):

thats basically what i was going to say :-D when doing problems like these, always look for things you can factor first. once you have factored what can be factored, look for like terms to cancel, then whatever is left is usually your answer

OpenStudy (anonymous):